# Manual-PA: Learning 3D Part Assembly from Instruction Diagrams

### v1

@@ 1 Introduction
With the rise of the DIY trend, nowadays it is the consumers who are increasingly assembling products like IKEA furniture [20, 2, 53, 54, 3], LEGO structures [46, 44], and exercise equipments [57, 42]. This paradigm shift can reduce the product cost, but often poses challenges for untrained consumers, as an incorrect assembly in any step may lead to irreparable outcomes. Thus, automating the assembly or providing active help in the process is often desired.
@@
Automated assembly faces the dual challenge of large combinatorial solution space, parameterized by the number of discrete parts in the assembly, and the need for a precise 6D pose estimation that correctly connects parts between each other. The limited number of feasible, stable assembly sequences makes this discrete-continuous optimization problem extremely challenging, even for advanced neural networks, especially as the number of parts increases [51]. Fortunately, most assembly tasks include language-agnostic instruction manuals, such as diagrams, delineating the assembly. While such manuals help shrink the solution search space, the problem of correctly matching the 2D manual diagrams with their corresponding 3D object parts and inferring their assembly pose and order remains a significant challenge for automated systems.
@@
This paper introduces a method to enable machines to learn how to assemble shapes by following a visual instruction manual, a task we call manual-guided 3D part assembly. As illustrated in Fig. 1, given a set of parts and a manual, our approach interprets the manual’s step-by-step diagrams to predict the precise 6D pose of each part, assembling them into the target shape. There are two lines of prior approaches that have similarities to our proposed task, namely methods that focus on the geometry of individual parts and inter-part relationships [22, 52, 55, 8, 56, 9], and methods such as MEPNet [46] that learns the assembly in specific conditions, such as LEGO shapes. The former rely on priors like peg-hole joints [24] or sequence generation [51] to reduce search complexity but are generative and may yield unstable assemblies due to occlusions [22]. The latter often simplifies the task by assuming that parts are provided step-by-step, e.g., LEGO manuals. However, furniture manuals, such as IKEA, may not have such information, and it is necessary to determine which parts are used at each step before assembly. Additionally, while LEGO tasks feature standardized “stud” joints, general furniture assembly is more complex and lacks such constraints.
@@
Based on the above observations, we identify two key challenges in our task. First, how to detect which part or parts are added at each step in the assembly diagrams. Given that these diagrams depict incremental assembly stages, the core challenge is to recognize newly added parts at each step. This requires aligning 2D assembly diagrams with the 3D parts to establish a correct assembly sequence. Second, how to effectively align the assembly process with the learned order. A straightforward approach is to directly use the learned sequence as the order for predicting poses, e.g., by adopting an auto-regressive prediction method similar to MEPNet [46]. However, this approach carries the risk of accumulating errors—if a mistake occurs in an earlier step, subsequent steps are likely to fail as well. On the other hand, completely disregarding the sequence forces the assembly model to implicitly find correspondences, which makes both learning and prediction significantly more difficult. Therefore, the inferred sequence should serve as a soft guidance for the assembly process, helping each part focus more on the relevant step diagram.
@@
To empirically validate the effectiveness of Manual-PA, we conduct experiments on the standard PartNet benchmark dataset [26], on the categories of assembling the Chair and Table classes. Our experiments and ablation studies clearly show that Manual-PA leads to state-of-the-art results on multiple metrics, including success rate, outperforming prior methods by a significant margin. Results further showcase the strong generalizability of our approach to real-world IKEA furniture assembly through experiments using the IKEA-Manual dataset [47].
@@
    We present experiments on the benchmark PartNet and IKEA-Manual datasets demonstrating state-of-the-art performances.
@@ 2 Related Works
involves reconstructing complete shapes from sets of 3D parts, a task originally assisted by leveraging semantic part information [6, 14, 49]. With the introduction of the PartNet dataset [26], recent works [52, 27, 8, 55, 9, 56] leverage geometric features. For instance, 3DHPA [9] uses part symmetry for a hierarchical assembly, SPAFormer [51] circumvents combinatorial explosion by generating a part order, and Joint-PA [24] introduces a peg-hole constraint approach to facilitate assembly. However, these methods are generative towards producing plausible final shapes without aiming for preciseness. Approaches like GPAT [23] and Image-PA [22] add guidance from point clouds or rendered images, respectively. Our work extends this direction by introducing guidance through instruction manuals, similar to how humans would do.
@@ 2.0.2 Permutation Learning
is often used in self-supervised methods for learning image representations [30, 39, 15, 4, 34]. However, in this work, we seek to align the 3D parts to instruction steps, for which we adopt a contrastive learning  [16, 13, 21, 38] using a novel cross-modal alignment setup.
@@
methods can be classified into instance-level [37, 35, 50], category-level [41, 45], and unseen [19, 48]. Our approach relates to the unseen object category as furniture typically comprises novel parts. When the 3D models of these parts are available, many methods [7, 33, 29, 25, 12] rely on feature matching followed by the PnP algorithm to infer pose. However, IKEA-style manuals are represented as line drawings making accurate feature extraction for matching to 3D parts challenging. Alternatively, some approaches [19, 48] use end-to-end models to directly predict the 6D pose. A common approach is to use CNOS [28] for object segmentation. However, in assembly diagrams, where multiple parts are connected and depicted together, SAM-based methods [17] often struggle with accurate segmentation. Our approach addresses these challenges using an end-to-end framework that uses a cross-attention scheme to enable each part to attend to its relevant region in the step diagram.
@@ 3.2 Feature Extraction
We first use off-the-shelf encoders to extract semantic latent features from the inputs. For the 3D part point clouds $\{\mathcal{P}_{i}\}_{i=1}^{N}$, we utilize a lightweight version of PointNet [36], similar to the one used in Image-PA [22], followed by a linear layer to map the features to a common dimension $D$, resulting in geometric part features ${\mathbf{f}^{\text{P}}}\in\mathbb{R}^{N\times D}$. For sequences of step diagrams $(\mathcal{I}_{j})_{j=1}^{N}$, given that the assembly instructions are provided step-by-step in an incremental manner, our primary focus is on the differences between consecutive steps. For any two consecutive diagram images $\mathcal{I}_{j}$ and $\mathcal{I}_{j+1}$, for $j=1,2,\ldots,N-1$, we compute the difference image given by $|\mathcal{I}_{j}-\mathcal{I}_{j+1}|$ that shows which new part is added to the thus-far assembled furniture. Next, this difference image is patchified into $K$ image patches, that are then fed to DINOv2 image encoder [5, 32] to extract features, followed by a linear layer to map these features to the same dimensionality $D$, resulting in semantic image features ${\mathbf{f}^{\text{I}}}\in\mathbb{R}^{N\times K\times D}$.
@@ 3.3 Manual-Guided Part Permutation Learning
Next, we obtain the permutation matrix $\mathbf{P}\in\{0,1\}^{N\times N}$ by solving a bipartite matching optimization problem using the Hungarian matching algorithm [18] with cost $\mathbf{C}=-\mathbf{S}$:
@@
where we minimize the cost, subject to constraints ensuring $\mathbf{P}$ is a permutation matrix, i.e., each row and column of $\mathbf{P}$ contains exactly one entry equal to $1$, with all other entries being 00. The final part order can be obtained by $\sigma=\operatorname{argmax}(\mathbf{P})$ where $\operatorname{argmax}$ is applied columnwise.
@@ 3.4 Manual-Guided Part Pose Estimation
Since the attention mechanism in the transformer is permutation-invariant, we follow Vaswani et al. [43] in using positional encoding to convey the order of both step diagrams and parts to the model. Formally, the positional encoding is defined as a function that maps a scalar to a sinusoidal encoding, $\phi:\mathbb{N}\rightarrow\mathbb{R}^{D}$:
@@
where $i=1,\dots,d/2$ indexes the dimension and $\tau_{p}$ is a temperature hyperparameter. By applying $\phi$ to encode the sequential order of step diagrams as they appear in the manual (i.e., using $x$ as a simple increasing sequence $1,2,\dots,N$), we obtain a positional encoding $\Phi\in\mathbb{R}^{N\times D}$, where ${\mathbf{p}^{\text{I}}}:=\Phi$. Given the permutation matrix $\mathbf{P}$ that corresponds to the order of the parts, we can compute the positional encoding for the parts as ${\mathbf{p}^{\text{P}}}=\mathbf{P}^{T}\Phi$. During training, we use the ground truth part order, whereas at inference time, we replace it with the predicted permutation matrix $\hat{\mathbf{P}}$, yielding $\hat{\mathbf{p}}^{\text{P}}=\hat{\mathbf{P}}^{T}\Phi$. These positional encodings are then added to the respective features before being fed into the transformer decoder and provide soft guidance on the sequence order.
@@
The modality interaction and fusion between step diagrams and parts occur through a stack of $L$ transformer [43] decoder layers. In each layer, the self-attention module receives the output from the previous layer (initially, the part features ${\mathbf{f}^{\text{P}}}+{\mathbf{p}^{\text{P}}}$ in the first layer) to perform part-to-part interactions. This is followed by a cross-attention module that injects and fuses the step diagram features ${\mathbf{f}^{\text{I}}}+{\mathbf{p}^{\text{I}}}$ with the part features processed by the self-attention module, enabling step-to-part interactions. Notably, for the step diagram features, we concatenate the features from all diagrams from the same manual along the patch dimension, resulting in a tensor of shape $\mathbb{R}^{NK\times D}$.
@@
Finally, the output $O_{\text{Dec}}$ from the transformer decoder is fed into a pose prediction head, which predicts the translation and rotation for each part, denoted as $(\hat{R}_{i},\hat{t}_{i})$. The rotation is represented as a unit quaternion $\hat{q}_{i}\in\mathbb{R}^{4}$. We initialize its bias to an identity quaternion $(1,0,0,0)$ and normalize the prediction to ensure a unit norm.
@@ 3.5 Training and Losses
where $B$ is the mini-batch size, $\tau$ is the temperature, and $\sigma(i)$ denotes the index for the $\sigma(i)$-th part that corresponds to the $j$-th step according to the ground truth part order $\sigma$. Here, mini-batches are constructed by sampling pairs of step-diagram and part features, $({\mathbf{f}^{\text{P}}}_{\!\!\sigma(i)},\textbf{g}^{\text{I}}_{j})$ for optimizing the InfoNCE loss [11, 31]. Note that we randomly sample one part from their geometric-equivalent group. In this setup, pairs $({\mathbf{f}^{\text{P}}}_{\!\!\sigma(i)},\textbf{g}^{\text{I}}_{i})$ are considered as positive, otherwise negative. This loss encourages positive pairs to have higher similarity while negative pairs have lower similarity, allowing the similarity matrix to be used for generating the permutation matrix via Hungarian matching, as discussed in Sec. 3.3.
@@
where CD denotes the chamfer distance measuring the similarity between unordered two point clouds [10], $i$ and $j$ index the parts within $\mathcal{G}$, and $\hat{R}$ and $\hat{t}$ denote the respective predicted rotation and translation. The optimal matching sequence is denoted as $\mathcal{M}$.
For the 3D part assembly task, following Li et al. [22], we supervise translation and rotation separately. We compute the $\ell_{2}$-distance for translation as follows:
@@ 4.1 Datasets
Following [52, 55, 8, 56, 9, 27, 51, 22], we use PartNet [26], a comprehensive 3D shape dataset with fine-grained, hierarchical part segmentation, for both training and evaluation. In our work, we adopt the train/validation/test split used by Li et al. [22], where shapes with more than 20 parts are filtered out, and we use the finest granularity, Level-3, of PartNet. We focus on the two largest furniture categories: Chair and Table, containing 40,148 and 21,517 parts, respectively. Besides that, to demonstrate the generalizability of our method, we also evaluate its zero-shot capabilities on the IKEA-Manual [47] dataset, which features real-world IKEA furniture at the part level that are aligned with IKEA manuals. For consistency with PartNet, we evaluate both the Chair category, with 351 and 143 parts, and the Table, with 143 parts.
@@
[…]
Chair Table Chair Table Chair Table
[…]
Manual-PA w/o Order Manual 3.0 3.6 79.07 74.03 34.13 37.71
Manual-PA (Ours) Manual 1.7 1.8 89.06 87.41 58.03 56.66
[…]
Manual-PA w/o Order Manual 12.8 8.9 38.36 42.01 1.754 10.53
Manual-PA (Ours) Manual 11.4 4.8 42.51 49.72 3.509 15.79
@@
Table 1: 3D part assembly results on the PartNet test split and Ikea-Manual. †: We re-trained the Image-PA model using diagrams (2D line drawing images) as the conditioning input instead of the original RGB images. The values in bold represent the best results, while underlined indicate the second.
4.2 Evaluation Metrics
To evaluate the effectiveness of 3D part assembly of different models, we employ three key metrics: Shape Chamfer Distance (SCD) [22], Part Accuracy (PA) [22], and Success Rate (SR) [23]. SCD quantifies the overall chamfer distance between predicted and ground truth shapes, providing a direct measure of assembly quality, scaled by a factor of $10^{3}$ for interpretability. PA assesses the correctness of individual part poses by determining whether the chamfer distance for a part falls below a threshold of $0.01$, reflecting the accuracy of each part’s alignment. Finally, SR evaluates whether all parts in an assembled shape are accurately predicted, offering a strict criterion for complete assembly success.
4.3 Main Results
4.3.1 Fully Supervised on PartNet.
In this setting, we train our model on PartNet for each category and test it on unseen shapes in the test split, which share similar patterns with the training set. As shown in Sec. 4.1, our method, Manual-PA, significantly outperforms all previous approaches in both the Chair and Table categories. Specifically, Manual-PA achieves superior performance compared to all other conditioned methods. Notably, when compared to Image-PA, we observe improvements of 26 and 17 part accuracy for the chair and table categories, respectively. This shows that for the 3D Part Assembly task, instruction manuals provide far better guidance than a single image or other conditioning inputs. “Manual-PA w/o Order” refers to our model without access to the part assembly order information, requiring it to learn this implicitly. Although the instruction manual provides strong visual guidance, performance improvements remain limited in this case, highlighting the challenges of implicit learning. By explicitly learning the correspondences between step diagrams and parts and applying positional encoding as soft guidance, we achieve PA improvements of approximately 10 and 13 in the chair and table categories, respectively. Note that all other methods, except for Image-PA and ours, are generative and follow the setting proposed by [52], where multiple assembly results are generated with varying Gaussian noise for the same set of parts. Only the most faithful assembly, as measured by Minimum Matching Distance (MMD) [1], is evaluated.
Figure 3: Comparison of average success rates (SR) across varying numbers of parts for different methods on PartNet chair test split. Number of chairs tested shown in background bar chart.
4.3.2 Number of Parts.
We analyze the assembly quality across varying numbers of parts on PartNet chair test split, as shown in Fig. 3. It is generally accepted that, assembling objects with more parts is more challenging due to the combinatorial explosive issue. However, our method consistently achieves a higher success rate across all part counts, demonstrating robust adaptability to different levels of assembly complexity. Notably, for shapes with more than 10 parts, 3DHPA and Image-PA exhibit near-zero success rates, whereas our method, Manual-PA, continues to produce competitive assembly results.
4.3.3 Zero-shot on IKEA-Manual.
We use the same model, pretrained on PartNet, to directly estimate the poses of furniture in IKEA-Manual, where all shapes are 3D models of real-world IKEA furniture, and each part corresponds to illustrations in their official assembly manuals. As shown in Sec. 4.1, our method, significantly outperforms strong competitors such as 3DHPA and Image-PA, demonstrating the strong generalizability of our approach to real world.
4.4 Ablation Studies
[…]
1 PE 4.3 72.22 22.50
2 PE ✓ 3.0 79.31 33.88
3 OneHot ✓ 3.0 79.07 34.13
4 OneHot ✓ ✓ 2.7 81.22 36.92
5 PE (Ours) ✓ ✓ 1.7 95.38 73.07
Table 2: Ablation results for various order designs on the PartNet chair test split. We use GT order for both training and inference.
[…]
Manual-PA (Ours) 0.789 1.7 89.06 58.03
 w/o batch-level sampling 0.410 2.3 56.93 15.68
 w/ Sinkhorn 0.774 1.7 88.25 56.26
 w/ GT Order 1.000 1.7 95.38 73.07
Table 3: Ablation results for various component designs in permutation learning on the PartNet chair test split. KT denotes the Kendall-tau coefficient that measure the ordinal correlation between predicted and GT order.
4.4.1 Effect of the Order for 3D Part Assembly.
We conduct experiments to explore various methods for incorporating order information in a 3D part assembly model, as shown in Sec. 4.4. First, by comparing lines 1 and 2, we observe that adding sequential information solely to part features improves assembly performance. We attribute this improvement to two factors: (1) distinguishing geometrically equivalent components, and (2) mitigating the combinatorial explosion problem, as claimed in SPAFormer [51]. Second, by comparing lines 2 and 5, where the latter provides both the step diagram and part features with correct correspondence, we achieve a significant performance boost. This indicates that explicitly establishing correspondence between steps and parts is beneficial, as learning these correspondences implicitly remains challenging for the model. Last, our empirical results show that, when position information is solely relevant to self-attention (lines 2 and 3), both PE and OneHot encodings yield similar performance. This observation aligns with prior transformer-based studies, which also utilize OneHot encodings [55, 9, 51]. However, in cases requiring cross-attention (line 4 and 5), PE outperforms OneHot encodings, highlighting its advantages in cross-attentive contexts.
4.4.2 Analysis of Permutation Learning Choices.
We conduct an ablation study for permutation learning, as shown in Sec. 4.4. First, we find that batch-level sampling is essential compared to applying contrastive learning individually within each manual book. This is primarily because contrastive learning benefits from a larger pool of negative samples. Second, while we employ the Sinkhorn [40] algorithm on the similarity matrix to enforce that rows and columns sum to one, empirical results indicate limited performance improvement. Last, by providing the model with the ground truth assembly order, we establish an upper performance bound for our assembly model.
Figure 4: Visualization of the cross attention scores between step diagrams (top) and a part (left, chair’s back) with PE at different positions. The final assembly results are displayed on the right.
Figure 5: Qualitative comparison of various 3D part assembly methods. Four examples are shown: chair (a) and table (b) from the PartNet dataset, and chair (c) and table (d) from the IKEA-Manual dataset.
4.5 Visualization and Analysis
4.5.1 Demonstration of Soft Guidance.
We manually add different positional encodings to the feature of chair’s back, as shown in Fig. 4. With PE(5), the model focuses on the final position, aligning well with the correct step diagram. However, when PE(2) is applied, drawing the model’s attention to the second step diagram, the model struggles to accurately predict the chair’s back pose. Notably, even with PE(0), the model manages to shift its attention toward the correct final step, demonstrating the fault tolerance and robustness of soft guidance: the model can still identify the correct match even when given an incorrect sequence.
4.5.2 Qualitative Comparison.
As shown in Fig. 5, our proposed Manual-PA produces the most faithful part pose prediction for the shapes illustrated in the step diagrams. For “Manual-PA w/o Order”, some parts do not align well with the assembly instructions, indicating the importance of providing precise correspondences between the step diagram and the parts for accurate assembly. In the case of Image-PA, the input consists only of the final image, making it challenging to locate the poses of parts that are occluded, such as the supporting rods beneath the table top in (b). 3DHPA lacks visual input and operates in a generative mode, relying solely on shape priors. Consequently, it generates shapes that are reasonable, as seen in (a), but there are noticeable discrepancies between its output and the ground truth shapes. For zero-shot results on the IKEA-Manual dataset, both 3DHPA and Image-PA exhibit degraded performance, while Manual-PA still produces faithful assembly results, demonstrating its superior generalization ability.
4.5.3 Cross Attention Map on Step Diagrams.
Figure 6: Visualization of cross-attention maps on step diagrams. The cross-attention represents the mean aggregation for each part across all step diagrams. Color red indicates a higher attention.
As shown in Fig. 6, we visualize the cross-attention maps between each part and the step diagrams. High attention values correspond to regions in the step diagram where the model pays more attention, which aligns with the spatial placement of parts during assembly. For instance, when the first blue board focuses on the back of the chair, it is positioned on the chair back, and when it attends to the area under the seat, it acts as a support between the legs. Comparing the two methods, we observe that without order as a soft guidance (represented by “Manual-PA w/o Order”), the attention regions are more dispersed, resulting in less accurate part placement. For example, the third cyan board does not focus entirely on the left chair leg without soft guidance, leading to a misaligned leg position. The absence of order as explicit guidance also results in incorrect part placement. For instance, the fourth purple board, which should be part of the seat, is instead assembled onto the chair back.
5 Conclusions and Future Work
We introduce a novel 3D part assembly framework that leverages diagrammatic manuals for guidance. Our proposed approach, Manual-PA, learns the sequential order of part assembly and uses it as a form of soft guidance for 3D part assembly. Manual-PA achieves superior performance on the PartNet dataset compared to existing methods and demonstrates strong generalizability to real-world IKEA furniture, as validated on the IKEA-Manual dataset.
Future work could aim to bridge the gap between synthetic and real-world IKEA manuals. First, the number of new parts introduced in each step can vary, requiring a more adaptable detection approach. Moreover, differing perspectives across steps necessitate robust viewpoint handling. The manuals also often include sub-module assemblies, resulting in a tree-like, non-linear assembly structure. Finally, while our current approach trains separate models for each furniture category, an ideal solution would be a unified model that generalizes across all kinds of furniture.

### v2

@@ 1 Introduction
With the rise of the DIY trend, nowadays it is the consumers who are increasingly assembling products like IKEA furniture [lee2021ikea, ben2021ikea, zhang2023aligning, zhang2024temporally, ben2024ikea], LEGO structures [wang2022translating, wan2024scanet], and exercise equipments [zhu2023multi, tian2022assemble]. This paradigm shift can reduce the product cost, but often poses challenges for untrained consumers, as an incorrect assembly in any step may lead to irreparable outcomes. Thus, automating the assembly or providing active help in the process is often desired.
@@
Automated assembly faces the dual challenge of large combinatorial solution space, parameterized by the number of discrete parts in the assembly, and the need for a precise 6D pose estimation that correctly connects parts between each other. The limited number of feasible, stable assembly sequences makes this discrete-continuous optimization problem extremely challenging, even for advanced neural networks, especially as the number of parts increases [xu2024spaformer]. Fortunately, most assembly tasks include language-agnostic instruction manuals, such as diagrams, delineating the assembly. While such manuals help shrink the solution search space, the problem of correctly matching the 2D manual diagrams with their corresponding 3D object parts and inferring their assembly pose and order remains a significant challenge for automated systems.
@@
This paper introduces a method to enable machines to learn how to assemble shapes by following a visual instruction manual, a task we call manual-guided 3D part assembly. As illustrated in Fig. 1, given a set of parts and a manual, our approach interprets the manual’s step-by-step diagrams to predict the precise 6D pose of each part, assembling them into the target shape. There are two lines of prior approaches that have similarities to our proposed task, namely methods that focus on the geometry of individual parts and inter-part relationships [li2020learning, zhan2020generative, zhang20223d, cheng2023score, zhang2024scalable, du2024generative], and methods such as MEPNet [wang2022translating] that learns the assembly in specific conditions, such as LEGO shapes. The former rely on priors like peg-hole joints [li2024category] or sequence generation [xu2024spaformer] to reduce search complexity but are generative and may yield unstable assemblies due to occlusions [li2020learning]. The latter often simplifies the task by assuming that parts are provided step-by-step, e.g., LEGO manuals. However, furniture manuals, such as IKEA, may not have such information, and it is necessary to determine which parts are used at each step before assembly. Additionally, while LEGO tasks feature standardized “stud” joints, general furniture assembly is more complex and lacks such constraints.
@@
Based on the above observations, we identify two key challenges in our task. First, how to detect which part or parts are added at each step in the assembly diagrams. Given that these diagrams depict incremental assembly stages, the core challenge is to recognize newly added parts at each step. This requires aligning 2D assembly diagrams with the 3D parts to establish a correct assembly sequence. Second, how to effectively align the assembly process with the learned order. A straightforward approach is to directly use the learned sequence as the order for predicting poses, e.g., by adopting an auto-regressive prediction method similar to MEPNet [wang2022translating]. However, this approach carries the risk of accumulating errors—if a mistake occurs in an earlier step, subsequent steps are likely to fail as well. On the other hand, completely disregarding the sequence forces the assembly model to implicitly find correspondences, which makes both learning and prediction significantly more difficult. Therefore, the inferred sequence should serve as a soft guidance for the assembly process, helping each part focus more on the relevant step diagram.
@@
To empirically validate the effectiveness of Manual-PA, we conduct experiments on the standard PartNet benchmark dataset [mo2019partnet], on the categories of assembling the chair, table and storage classes. Our experiments and ablation studies clearly show that Manual-PA leads to state-of-the-art results on multiple metrics, including success rate, outperforming prior methods by a significant margin. Results further showcase the strong generalizability of our approach to real-world IKEA furniture assembly through experiments using the IKEA-Manual dataset [wang2022ikea].
@@
    We present experiments on the benchmark PartNet and IKEA-Manual datasets demonstrating state-of-the-art performances and cross-category generalization.
@@ 2 Related Works
involves reconstructing complete shapes from sets of 3D parts, a task originally assisted by leveraging semantic part information [funkhouser2004modeling, sung2017complementme, yin2020coalesce, chaudhuri2011probabilistic, kalogerakis2012probabilistic, wu2020pq]. With the introduction of the PartNet dataset [mo2019partnet], recent works [zhan2020generative, narayan2022rgl, cheng2023score, zhang20223d, du2024generative, zhang2024scalable] leverage geometric features. For instance, 3DHPA [du2024generative] uses part symmetry for a hierarchical assembly, SPAFormer [xu2024spaformer] circumvents combinatorial explosion by generating a part order, and Joint-PA [li2024category] introduces a peg-hole constraint to facilitate assembly. However, these methods are generative towards producing plausible final shapes without aiming for preciseness. Approaches like GPAT [li2023rearrangement] and Image-PA [li2020learning] add guidance from point clouds or rendered images, respectively. Our work extends this direction by introducing guidance through instruction manuals, similar to how humans would do.
@@ 2.0.2 Permutation Learning
is often used in self-supervised methods for learning image representations [noroozi2016unsupervised, santa2017deeppermnet, kim2018learning, carlucci2019domain, pang2020solving]. However, in this work, we seek to align the 3D parts to instruction steps, for which we adopt a contrastive learning  [kim2021vilt, jia2021scaling, li2021align, radford2021learning] using a novel cross-modal alignment setup.
@@
methods can be classified into instance-level [rad2017bb8, peng2019pvnet, xiang2017posecnn], category-level [tian2020shape, wang2019normalized], and unseen [labbe2022megapose, wen2024foundationpose]. Our approach relates to the unseen object category as furniture typically comprises novel parts. When the 3D models of these parts are available, many methods [chen2023zeropose, ornek2023foundpose, nguyen2024gigapose, lin2024sam, huang2024matchu] rely on feature matching followed by the PnP algorithm to infer pose. However, IKEA-style manuals are represented as line drawings making accurate feature extraction for matching to 3D parts challenging. Alternatively, some approaches [labbe2022megapose, wen2024foundationpose] use end-to-end models to directly predict the 6D pose. A common approach is to use CNOS [nguyen2023cnos] for object segmentation. However, in assembly diagrams, where multiple parts are connected and depicted together, SAM-based methods [kirillov2023segment] often struggle with accurate segmentation. Our approach addresses these challenges using an end-to-end framework that uses a cross-attention scheme to enable each part to attend to its relevant region in the step diagram.
@@ 3.2 Feature Extraction
We first use off-the-shelf encoders to extract semantic latent features from the inputs. For the 3D part point clouds $\{\mathcal{P}_{i}\}_{i=1}^{N}$, we utilize a lightweight version of PointNet [qi2017pointnet], similar to the one used in Image-PA [li2020learning], followed by a linear layer to map the features to a common dimension $D$, resulting in geometric part features ${\mathbf{f}^{\text{P}}}\in\mathbb{R}^{N\times D}$. For sequences of step diagrams $(\mathcal{I}_{j})_{j=1}^{N}$, given that the assembly instructions are provided step-by-step in an incremental manner, our primary focus is on the differences between consecutive steps. For any two consecutive diagram images $\mathcal{I}_{j}$ and $\mathcal{I}_{j+1}$, for $j=1,2,\ldots,N-1$, we compute the difference image given by $|\mathcal{I}_{j}-\mathcal{I}_{j+1}|$ that shows which new part is added to the thus-far assembled furniture. Next, this difference image is patchified into $K$ image patches, that are then fed to DINOv2 image encoder [caron2021emerging, oquab2023dinov2] to extract features, followed by a linear layer to map these features to the same dimensionality $D$, resulting in semantic image features ${\mathbf{f}^{\text{I}}}\in\mathbb{R}^{N\times K\times D}$.
@@ 3.3 Manual-Guided Part Permutation Learning
Next, we obtain the permutation matrix $\mathbf{P}\in\{0,1\}^{N\times N}$ by solving a bipartite matching optimization problem using the Hungarian matching algorithm [kuhn1955hungarian] with cost $\mathbf{C}=-\mathbf{S}$:
@@
where we minimize the cost, subject to constraints ensuring $\mathbf{P}$ is a permutation matrix, i.e., each row and column of $\mathbf{P}$ contains exactly one entry equal to $1$, with all other entries being 0. The final part order can be obtained by $\sigma=\operatorname{argmax}(\mathbf{P})$ where $\operatorname{argmax}$ is applied columnwise.
@@ 3.4 Manual-Guided Part Pose Estimation
Since the attention mechanism in the transformer is permutation-invariant, we follow vaswani2017attention [vaswani2017attention] in using positional encoding to convey the order of both step diagrams and parts to the model. Formally, the positional encoding is defined as a function that maps a scalar to a sinusoidal encoding, $\phi:\mathbb{N}\rightarrow\mathbb{R}^{D}$:
@@
where $i=1,\dots,d/2$ indexes the dimension and $\tau_{p}$ is a temperature hyperparameter. By applying $\phi$ to encode the sequential order of step diagrams as they appear in the manual (i.e., using $x$ as a simple increasing sequence $1,2,\dots,N$), we obtain a positional encoding $\Phi\in\mathbb{R}^{N\times D}$, where ${\mathbf{p}^{\text{I}}}:=\Phi$. Given the permutation matrix $\mathbf{P}$ that corresponds to the order of the parts, we can compute the positional encoding for the parts as ${\mathbf{p}^{\text{P}}}=\mathbf{P}^{T}\Phi$. During training, we use the ground truth part order, whereas at inference time, we replace it with the predicted permutation matrix $\hat{\mathbf{P}}$, yielding $\hat{\mathbf{p}}^{\text{P}}=\hat{\mathbf{P}}^{T}\Phi$. These positional encodings are then added to the respective features before being fed into the transformer decoder and provide soft guidance on the sequence order. We adopt RoPE [su2024roformer] for better performance.
@@
The modality interaction and fusion between step diagrams and parts occur through a stack of $L$ transformer [vaswani2017attention] decoder layers. In each layer, the self-attention module receives the output from the previous layer (initially, the part features ${\mathbf{f}^{\text{P}}}+{\mathbf{p}^{\text{P}}}$ in the first layer) to perform part-to-part interactions. This is followed by a cross-attention module that injects and fuses the step diagram features ${\mathbf{f}^{\text{I}}}+{\mathbf{p}^{\text{I}}}$ with the part features processed by the self-attention module, enabling step-to-part interactions. Notably, for the step diagram features, we concatenate the features from all diagrams from the same manual along the patch dimension, resulting in a tensor of shape $\mathbb{R}^{NK\times D}$.
@@
Finally, the output $O_{\text{Dec}}$ from the transformer decoder is fed into a pose prediction head, which predicts the translation and rotation for each part, denoted as $(\hat{R}_{i},\hat{t}_{i})$. In our implementation, the rotation matrix $\hat{R}_{i}$ is represented as a unit quaternion $\hat{q}_{i}\in\mathbb{R}^{4}$. We initialize its bias to an identity quaternion $(1,0,0,0)$ and normalize the prediction to ensure a unit norm.
@@ 3.5 Training and Losses
where $B$ is the mini-batch size, $\tau$ is the temperature, and $\sigma(i)$ denotes the index for the $\sigma(i)$-th part that corresponds to the $j$-th step according to the ground truth part order $\sigma$. Here, mini-batches are constructed by sampling pairs of step-diagram and part features, $({\mathbf{f}^{\text{P}}}_{\!\!\sigma(i)},\textbf{g}^{\text{I}}_{j})$ for optimizing the InfoNCE loss [hadsell2006dimensionality, oord2018representation]. Note that we randomly sample one part from their geometric-equivalent group. In this setup, pairs $({\mathbf{f}^{\text{P}}}_{\!\!\sigma(i)},\textbf{g}^{\text{I}}_{i})$ are considered as positive, otherwise negative. This loss encourages positive pairs to have higher similarity while negative pairs have lower similarity, allowing the similarity matrix to be used for generating the permutation matrix via Hungarian matching, as discussed in Sec. 3.3.
@@
where CD denotes the chamfer distance measuring the similarity between unordered two point clouds [fan2017point], $i$ and $j$ index the parts within $\mathcal{G}$, and $\hat{R}$ and $\hat{t}$ denote the respective predicted rotation and translation. The optimal matching sequence is denoted as $\mathcal{M}$.
For the 3D part assembly task, following li2020learning [li2020learning], we supervise translation and rotation separately. We compute the $\ell_{2}$-distance for translation as follows:
@@ 4.1 Datasets
Following [zhan2020generative, zhang20223d, cheng2023score, zhang2024scalable, du2024generative, narayan2022rgl, xu2024spaformer, li2020learning], we use PartNet [mo2019partnet], a comprehensive 3D shape dataset with fine-grained, hierarchical part segmentation, for both training and evaluation. In our work, we adopt the train/validation/test split used by li2020learning [li2020learning], where shapes with more than 20 parts are filtered out, and we use the finest granularity, Level-3, of PartNet. We focus on the three largest furniture categories: chair, table and storage, containing 40,148, 21,517 and 9,258 parts, respectively. Besides that, to demonstrate the generalizability of our method, we also evaluate its zero-shot capabilities on the IKEA-Manual [wang2022ikea] dataset, which features real-world IKEA furniture at the part level that are aligned with IKEA manuals. For consistency with PartNet, we evaluate both chair category with 351 parts, and table, with 143 parts. Storage is excluded due to only 3 samples.
@@
Table 1: 3D part assembly results on the PartNet test split and Ikea-Manual. †: We re-trained the Image-PA model using diagrams (2D line drawing images) as the conditioning input instead of the original RGB images. The values in bold represent the best results, while underlined indicate the second. ‡: Average over five runs.
\__nicematrix_patch_booktabs:\__nicematrix_revert_colortbl: